OpenStudy (anonymous):

MEDAL!!!!! Points A, B, and C are collinear. Point B is between A and C. Solve for x. 11) AC = 3x+3, AB = -1+2x, and BC = 11. Find x. SHOW WORK

OpenStudy (anonymous):

is this on that paper too?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

damn does it say point B is half way between A and C

OpenStudy (anonymous):

oh nvm, it is the same as the last one

OpenStudy (anonymous):

\[AC=AB+BC\] \[3x+3=-1+2x+11\] solve for \(x\)

OpenStudy (anonymous):

lazy math teacher with dull, dumb algorithmically generated questions from kutasoftware ought to be ashamed

OpenStudy (anonymous):

you need help solving that for \(x\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

combine like terms, get \[3x+3=2x+10\] subtract \(2x\) get \[x+3=10\] then subtract \(3\) to get \(x=7\)

OpenStudy (anonymous):

then plug that back in to \(3x+3\) to get \[3\times 7+3=24\]

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

yw